OpenStudy (anonymous):

you are dealt 5 cards out of a 52 card deck. what is the probability you are dealt 2 kings and 2 sevens?

OpenStudy (raden):

you tell that dealt 5 cards, but ur question dealt 2 (K) and 2 (7)'s, where is the rest one ? :)

OpenStudy (anonymous):

it could be any of the remaining cards

OpenStudy (anonymous):

i keep getting (4 choose 2)(4 choose 2)(44 choose 1)/(52 choose 2)

OpenStudy (sirm3d):

two kings, two sevens, and a card other can K or 7 \[\frac{\displaystyle \binom{4}{2}\binom{4}{2}{\binom{44}{1}}}{\displaystyle \binom{52}{5}}\]

OpenStudy (sirm3d):

why (52 choose 2). shouldn't you choose 5 cards among 52?

OpenStudy (anonymous):

yep sorry thats what i ment

OpenStudy (anonymous):

this is what i used on wolfram http://www.wolframalpha.com/input/?i=probability+2+kings+and+2+queens

OpenStudy (anonymous):

so it keeps showing that the numerator should be 612

OpenStudy (shubhamsrg):

KKQQE <- 4/52 * 3/51 * 4/50 * 3/49 * 44/48 KKQEQ <- 4/52 * 3/51 * 4/50 * 44/49 * 3/48 So for every permutation you'll get (4*4*3*3*44)/(52*52*50*49*48) total no. of permutations = 5!/2! 2! = 30 Even my logic yield the same ans as what we obtained here and not wolfram's. hmm

OpenStudy (anonymous):

alright see i have gone though this many times. maybe wolfram is taking something into consideration that im not but its alright. I think im gunna just go with what i originally had and it should be ok. Thank you for your help!

OpenStudy (zarkon):

wolfram appears to be doing some crazy stuff